Suppose your portfolio invest in two securities A and B. The proportional bid ask spread for two securities are independent and normally distributed estimated as mean -0.5% and Sigma -0.1% for first security and mean -0.4% and Sigma -0.05% for second security. The current value of your portfolio is worth $20,000,000. The investment weights on each security are 60% and 40%.

Please estimate the liquidity risk of your portfolio with 99% confidence level.
Answer #1

Ans-

for stock A

Given,

mean=-0.5%

sigma=-0.1%

portfolio value=12000000

Now Liquidity risk is

VaR= (sigma - mean) * 1.65 * portfolio value=(-0.5%+0.1%)*1.65*12000000=-98980.2

FOr B

VaR=(sigma - mean) * 1.65 * portfolio value=(-0.05%+0.4%)*1.65*8000000=46200